OpenStudy (faf3):

Solve -3e^(x+1)-21= -27

OpenStudy (anonymous):

e^(x+1)=2

OpenStudy (anonymous):

x+1=ln2

OpenStudy (anonymous):

x=ln2-1

OpenStudy (anonymous):

x=-0.3

OpenStudy (anonymous):

(x+4)(x-2)=7

OpenStudy (anonymous):

x^2+2x-15

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

(x+5)(x-3)=0

OpenStudy (anonymous):

x=-5 or 3

OpenStudy (faf3):

thank you. ln(5x-3)-lnx= ln2

OpenStudy (anonymous):

5x-3=2x

OpenStudy (anonymous):

x=1

OpenStudy (faf3):

\[\log_{3}(x-24)+\log_{3}x= 4 \]

OpenStudy (anonymous):

x(x-24)=81

OpenStudy (anonymous):

x^2-24x-81=0

OpenStudy (anonymous):

x^2-27x+3x-81=0

OpenStudy (anonymous):

x(x-27)+3(x-27)=0

OpenStudy (anonymous):

(x+3)(x-27)=0

OpenStudy (anonymous):

x=-3 or 27

OpenStudy (faf3):

log(x+4)=log(x-2)+1

OpenStudy (anonymous):

(x+4)/(x-2)= 10

OpenStudy (anonymous):

x+4=10x-20

OpenStudy (anonymous):

x=24/9=8/3

OpenStudy (faf3):

Al wants to invest $50,000. His options are either to invest his money into an account paying 1.75% a.p.r. compounded quarterly for 3 years or he can put his money into an account paying 1.7% a.p.r. compounded continuously for 3 years. Which option should he choose? How much extra money would reproduced in the more profitable account?